Andy is twice as old as kate. In years, their ages will total 60. how old is each now?

Answer by avani(25)

LET ANDY AGE =2X and kATE AGE = X
IN YEARS, THE ADDITION OF THEIR AGE =60
SO,
2X+X=60
3X=60
X=60/3=20
THUS KATE IS 20 YEARS OLD AND ANDY IS 2(20)=40YEARS OLD